# Fusion rule in conformal field theories and topological orders: A unified view of correspondence and (fractional) supersymmetry and their relation to topological holography

### v5

@@ Abstract
The algebraic or ring structure of anyons, called the fusion rule, is one of the most fundamental research interests in contemporary studies on topological orders (TOs) and the corresponding conformal field theories (CFTs). Recently, the algebraic structure realized as generalized symmetry, including non-invertible and categorical symmetry, captured attention in the fields. Such non-abelian anyonic objects appear in a bulk CFT or chiral CFT (CCFT), but it has been known that the construction of a CCFT contains theoretical difficulties in general. In this work, we study the structure of the fusion rules in $Z_{N}$ symmetric chiral and bulk conformal field theories and the corresponding TOs. We propose a nontrivial expression of subalgebra structure in the fusion rule of a bulk CFT. We name this subalgebra “bulk semion". This corresponds to the fusion rule of the CCFT and categorical symmetry of the TO or symmetry topological field theory (SymTFT). This gives a bulk-edge correspondence based on the symmetry analysis and corresponds to an anyon algebraic expression of topological holography in the recent literature. The recent topological holography is expected to apply to systems in general space-time dimensions. Moreover, we present a concise way of unifying duality (or fractional supersymmetry), generalized or categorical symmetry, and Lagrangian subalgebra. Our method is potentially useful to formulate and study general TOs, fundamentally only from the data of bulk CFTs or vice versa, and gives a clue in understanding CCFT (or ancillary CFT more generally).
@@
Recently, categorical symmetry or symmetry topological field theory (SymTFT) [19, 20, 21, 22] has captured attention as a symmetry-based description of TOs. For example, the double semion algebra in the $1+1$ dimensional Majorana-Ising CFT or in the $2+1$ dimensional $Z_{2}$ TO has been studied. Historically, the correspondence between the double semion algebra and the Majorana-Ising CFT appeared in the study of fermionic string theories[23]. However, generalization to a $Z_{N}$ symmetric model has not been studied except for several works[24, 25, 26, 27], regardless of the fundamental importance for the understanding of properties of anyons. It is worth stressing that the resultant $Z_{N}$ extended chiral CFT (or fractional supersymmetric string theory in the context of $Z_{N}$ nonanomalous models as in [28]) is outside of existing modular tensor categories (MTCs) and has been lacking general mathematical frameworks[25, 29, 30]. In other words, surprisingly, even the fusion algebraic data of anyons corresponding to fractional quantum Hall (FQH) states in condensed matter has not been studied sufficiently. It is also worth noting that an FQH state is regarded as a typical example of symmetry-enriched topological orders in more recent literature.
@@
In this work, we propose a general correspondence between a bulk CFT (product of chiral and antichiral CFTs) or spherical fusion category (SFC) and a chiral CFT (CCFT) or TO (or SymTFT) at the level of the fusion rule of anyonic objects. Starting from a $Z_{N}$ symmetric bulk CFT and its fusion algebra, we show a nontrivial construction of subalgebra based on its symmetry charge (or quantum anomaly) and parity (or particle counting) of the objects. We name this subalgebra “bulk semion algebra," corresponding to the algebraic (or categorical) structure of the TOs. The CFT/TQFT results in the correspondence between the bulk semion algebra and the fusion rule of CCFT. Our formalism provides the necessary algebraic data formulating “topological holography" [31, 32, 33, 34, 35, 36, 37], symmetry-based representation of the CFT/TQFT [38, 3, 39, 40, 41, 42] (However, we note that the term “topological holography" appeared earlier to notify the system both topological and holographic in [43, 44, 45, 46] in a different context. Similar terminological issues have been summarized in the introduction of [47]. We also note that the almost same term has been used in optics with different meanings[48, 49]). Moreover, our formalism is much more concise and general, with intuitive algebraic expressions. The main proposal of the present manuscript can be summarized as the following schematic picture,
@@
where $\rightarrow$ represents the $Z_{N}$ extension and $\Rightarrow$ represents the bulk semionization. The arrow $\rightarrow$ transforms a $Z_{N}$ symmetric bosonic bulk CFT to a $Z_{N}$ extended bulk CFT analogous to a quark-hadron like model, or a fractional supersymmetric model. In abstract algebra, the topological holography is just a combination of the group extension (or the operation taking a tensor product between a group ring and SFC) and the operation taking a subalgebra in our settings, and this can be expressed the combination of the two arrows. We also note this kind of reduction from a bulk CFT($\sim\text{CCFT}\times\overline{\text{CCFT}}$) to a CCFT can be interpreted as a massive renormalization group flow[50, 51, 52] which has a close connection to the Li-Haldane conjecture in condensed matter physics[53, 54, 55] and Moore-Seiberg data in high energy physics[56, 57]. In $2+1$ dimensional TOs or TQFTs, the above arrow $\Rightarrow$ corresponds to the gluing process of chiral and antichiral edges (See also the coresponding discussions in [58, 54, 26]). We also remind that the TOs with two edges, or those in the cylinder geometry, correspond to modular partition functions. Historically, the significance of the cylinder geometry and its correspondence to the bulk CFT have been established in earlier literature[59, 60].
The rest of the manuscript is organized as follows. Sec. II is the main part of this work, and we propose a construction of bulk semion algebra which corresponds to the fusion rule of anyonic objects in $2+1$ dimensional $Z_{N}$ TOs or the SymTFT. In Sec.III, we discuss the application of our method to construct correspondence between TOs and the CCFTs. Moreover, we report a new representation of the bulk CFT which can be related to the (fractional) supersymmetry or related even-odd problem on the lattice model. In Sec. IV, we show the precise relationship between our algebraic formalism and existing (and conjectural) category theories. This section provides an overview of the understanding of TOs with respect to anyon algebra. In Sec.V, we make a concluding remark by showing a figure of the correspondence between CFTs and TOs, so-called topological holography, potentially applicable to systems in general space-time dimensions.
@@ II Bulk semionization
In this section, we present the main results of this work, the $Z_{N}$ symmetric bulk semionization of anomaly-free $Z_{N}$ two-dimensional conformal field theories (summarized as the construction of a SymTFT $\mathbf{S}^{(l)}$ from an SFC $\mathbf{F}^{(l)}$). This provides an algebraic way to relate topological symmetry or SFC and categorical symmetry or SymTFT. In the following discussion, we use the same notations as in the author’s previous works[61, 27] and note the reviews [62, 63] for earlier literature on the simple current extension.
@@
where $\times$" represents the fusion of objects, $h$ is the conformal dimension, and we do not distinguish the label and the operator itself to simplify the notations.
The anomaly-freeness of this $Z_{N}$ theory can be defined as the condition $Q_{J}(J^{n})=0$ for all n=0,1,,N1n=0,1,...,N-1 where $J^{n}$ is the $n$ times multiplication of $J$. In other words, the $Z_{N}$ symmetry itself is not charged (or not mutually twisted) in the anomaly-free theory. This condition can be satisfied under the following conditions,
@@
The models satisfying these conditions have been known as discrete torsion or integer simple current extension of CFTs [64, 65, 66, 67]. For example, this condition can be realized in the $SU(N)_{N}$ Wess-Zumino-Witten (WZW) model[68, 69]. More recently, its relation to $SU(N)$ Haldane conjecture[70, 71, 72] or Lieb-Schultz-Mattis theorem has been revisited in [73, 61]. For the readers interested in the anomaly analysis of the $Z_{N}$ symmetric models and their renormalization group (RG) flows, we note several recent works [74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84] and earlier related works[85, 86, 87, 88, 89] and note recent works on noninvertible symmetry [90, 91, 92, 93, 94, 95, 96, 97, 47]. The anomaly-free $Z_{N}$ models play a central role in constructing $Z_{N}$ TOs or FQH states [98, 99, 24, 100, 101, 61, 102, 103] with close analogy and connection to quark-hadron models or fractional supersymmetric models.
@@
In other words, the algebraic data of the SFC are constrained by the algebraic data of the MTC, and one can consider that a class of SFCs is determined by the combination of the Verlinde formula [5, 6, 7] and Moore-Seiberg data [56, 57]. This can be useful for the readers to understand the SFC from the knowledge of MTC or CCFT.
The simple current extended bulk CFT $\mathbf{F}$ [64, 65, 66, 67] can be obtained by introducing the parity shift operation [117] (or recursive multiplications of $J$ and $\overline{J}$) to the bosonic theory $\mathbf{B}$. Alternatively, the simple current extension of CCFT can be realized by decomposing the $Z_{N}$ invariant sector as $\theta_{a,0}=\sum_{p=0}^{N-1}\phi_{a,p}/\sqrt{N}$ where the extended object $\phi_{a,p}$ is the $Z_{N}$ generalization of chiral semion[23, 118, 61, 27]. Because of the introduction of this extended object, the extended CCFT does not correspond to an MTC. Under the operator-state correspondence[27], the bulk operators can be specified as,
@@
where $Q$ is the $Z_{N}$ charge of the sectors.
@@
As one can see from the above expressions, the summation has been taken corresponding to the addition of the anomaly-free particle $J\overline{J}^{N-1}$, and this can be interpreted as a natural extension of the boson condensation. Whereas the usual anyon condensation results in ring isomorphism or ring homomorphism, the above condensation results in a subalgebra of $\mathbf{F}$. Because of the anomaly-free condition, the total $Z_{N}$ parity and charge are well-defined after the bulk semionization. Hence, one can relate the bosonic or $Z_{N}$ orbifolded modular invariants to the Lagrangian subalgebra of $\mathbf{S}$ [136, 26, 61] based on them by taking the parity or charge 0 sector. To emphasize the role of subalgebra as anyon condensation, we use Lagrangian subalgebra instead of “Lagrangian algebra" in the existing literature.
@@ II.1 Lagrangian (sub)algebra
One can check their consistencies by applying the fusion rule of $\mathbf{F}^{(0)}$ and changing the basis, because $\mathbf{S}^{(0)}$ is a subalgebra of $\mathbf{F}^{(0)}$. For example, one can check the consistency of the new identity operator $\mathbf{I}$ as follows,
@@ III Topological holography and fractional supersymmetry
where we have used the same notations in Sec.II. This is the algebraic representation of topological holography[32, 33, 34, 35, 36, 37] relating a $2+1$ dimensional TO denoted by $\{\Psi\}$ and a CCFT denoted by $\{\phi\}$. As an earlier literature in mathematics, we note [39] with detailed interpretations of the Moore-Seiberg data[56, 57]. We also note that the $Z_{N}$ symmetric Cardy states in [153, 154, 155, 156, 61, 34] which has a close connection to the bulk topological degeneracies can be labeled by the chiral semions because of the close connection to CCFTs and boundary CFTs (BCFTs)[51, 52, 157, 158, 159, 160] or the Li-Haldane conjecture in condensed matter[53, 54, 55].

### v7

@@ Abstract
The algebraic or ring structure of anyons, called the fusion rule, is one of the most fundamental research interests in contemporary studies on topological orders (TOs) and the corresponding conformal field theories (CFTs). Recently, the algebraic structure realized as generalized symmetry, including non-invertible and categorical symmetry, captured attention in the fields. Such non-abelian anyonic objects appear in a bulk CFT or chiral CFT (CCFT), but it has been known that the construction of a CCFT contains theoretical difficulties in general. In this work, we propose the fusion rules in $Z_{N}$ extended chiral and bulk conformal field theories and the corresponding TOs. We explicitly construct a nontrivial expression of subalgebra structure in the fusion rule of a bulk CFT. We name this subalgebra “bulk semion". This corresponds to the fusion rule of the CCFT and categorical symmetry of the TO or $Z_{N}$ graded symmetry topological field theory (SymTFT). This gives a bulk-edge correspondence based on the symmetry analysis and corresponds to an anyon algebraic expression of topological holography in the recent literature. The recent topological holography is expected to apply to systems in general space-time dimensions. Moreover, we present a concise way of unifying duality (or fractional supersymmetry), generalized or categorical symmetry, and Lagrangian subalgebra. Our method is potentially useful to formulate and study general TOs, fundamentally only from the data of bulk CFTs or vice versa, and gives a clue in understanding CCFT (or ancillary CFT more generally).
@@
Recently, categorical symmetry or symmetry topological field theory (SymTFT) [19, 20, 21, 22] has captured attention as a symmetry-based description of TOs. For example, the double semion algebra in the $1+1$ dimensional Majorana-Ising CFT or in the $2+1$ dimensional $Z_{2}$ TO has been studied. Historically, the correspondence between the double semion algebra and the Majorana-Ising CFT appeared in the study of fermionic string theories[23]. However, generalization to a $Z_{N}$ symmetric model has not been studied except for several works[24, 25, 26, 27], regardless of the fundamental importance for the understanding of properties of anyons. It is worth stressing that the resultant $Z_{N}$ extended chiral CFT (or fractional supersymmetric string theory in the context of $Z_{N}$ nonanomalous models as in [28]) is outside of existing modular tensor categories (MTCs) and has been lacking general mathematical frameworks[25, 29, 30]. In other words, surprisingly, even the fusion algebraic data of anyons corresponding to fractional quantum Hall (FQH) states, $Z_{N}$ graded SymTFTs, have not been studied sufficiently (Related unusual properties have been clarified more recently in [27]). It is also worth noting that an FQH state is regarded as a typical example of symmetry-enriched topological orders in more recent literature. More phenomenologically,the chiral correlation functions of the anomaly-free $Z_{N}$ particle, called (half) integer spin simple current appearing in the studies on discrete torsion and extended models[31, 32, 33, 34, 35, 36, 37], define $Z_{N}$ generalization of the wavefunctions of the corresponding FQH states or TQFTs through the bulk-edge correspondence or CFT/TQFT correspondence[38, 3]. The corresponding fusion rules are nothing but $Z_{N}$ graded SymTFTs in a modern technical term, and this unusual structure outside of bosonic theory has been studied several times, for example in [39, 40, 24].
@@
In this work, we propose a new $Z_{N}$ extended version of general correspondence between a bulk CFT (product of chiral and antichiral CFTs) or spherical fusion category (SFC) and a chiral CFT (CCFT) or TO (or SymTFT) at the level of the fusion rule of anyonic objects. Starting from a $Z_{N}$ symmetric bulk CFT and its fusion algebra, we show a nontrivial construction of subalgebra, $Z_{N}$ graded SymTFT denoted as $\mathbf{S}$, based on its symmetry charge (or quantum anomaly) and parity (or particle counting) of the objects. We name this subalgebra “bulk semion algebra," corresponding to the algebraic (or categorical) structure of the TOs. The CFT/TQFT results in the correspondence between the bulk semion algebra and the fusion rule of CCFT. Our formalism provides the necessary algebraic data formulating “topological holography" [41, 42, 43, 44, 45, 46, 47], symmetry-based representation of the CFT/TQFT [48, 3, 49, 50, 51, 52] (However, we note that the term “topological holography" appeared earlier to notify the system both topological and holographic in [53, 54, 55, 56] in a different context. Similar terminological issues have been summarized in the introduction of [57]. We also note that the almost same term has been used in optics with different meanings[58, 59]). Moreover, our formalism is much more concise and general, with intuitive and rigorous algebraic expressions with a concrete identification with modular partition functions, Eq.(13), and analysis on $Z_{N}$ anomaly or fractional supercharge defined by Eq.(2) that are absent in references in category theory. The corresponding category theories are still under development, and our explicit method from the established techniques in abstract algebra (or linear algebra) will provide a clue in constructing such extended category theories in a coherent way.
The main proposal of the present manuscript can be summarized as the following schematic picture,
@@
where $\rightarrow$ represents the $Z_{N}$ extension and $\Rightarrow$ represents the bulk semionization. The arrow $\rightarrow$ transforms a $Z_{N}$ symmetric bosonic bulk CFT to a $Z_{N}$ extended bulk CFT analogous to a quark-hadron-like model, or a fractional supersymmetric model. This operation itself has been studied in category theory[60, 61], but the resultant partition functions, Eq. (13), and their anomaly-free conditions have not been studied in this context to our knowledge. In abstract algebra, the topological holography is just a combination of the group extension (or the operation taking a tensor product between a group ring and SFC) and the operation taking a subalgebra in our settings, and this can be expressed as the combination of the two arrows. This simplification by abstract algebra provides a powerful theoretical framework for exact computations of fusion coefficients in the unexplored topological orders, $Z_{N}$ graded SymTFTs. We also note that this kind of reduction from a bulk CFT($\sim\text{CCFT}\times\overline{\text{CCFT}}$) to a CCFT can be interpreted as a massive renormalization group flow[62, 63, 64] which has a close connection to the Li-Haldane conjecture in condensed matter physics[65, 66, 67] and Moore-Seiberg data in high energy physics[68, 69]. In $2+1$ dimensional TOs or TQFTs, the above arrow $\Rightarrow$ corresponds to the gluing process of chiral and antichiral edges (See also the corresponding discussions in [70, 66, 26]). We also remind that the TOs with two edges, or those in the cylinder geometry, correspond to modular partition functions. Historically, the significance of the cylinder geometry and its correspondence to the bulk CFT have been established in earlier literature[71, 72].
The rest of the manuscript is organized as follows. Sec. II is the main part of this work, and we propose a construction of the bulk semion algebra or $Z_{N}$ graded SymTFT $\mathbf{S}$ which corresponds to the fusion rule of anyonic objects in $2+1$ dimensional $Z_{N}$ TOs along with discussions on the less familiar (or new) modular partition functions, Eq.(13). In Sec.III, we discuss the application of our method to construct correspondence between TOs and the CCFTs. Consistency between our method and existing chiral algebra is checked in the Majorana CFT. Moreover, we report a new representation of the bulk CFT which can be related to the (fractional) supersymmetry or related even-odd problem on the lattice model along with the unusual degenerate fusion rule. In Sec. IV, we show the precise relationship between our algebraic formalism and existing (and conjectural) category theories. This section provides an overview of the understanding of TOs with respect to anyon algebra. In Sec.V, we make a concluding remark by showing a figure of the correspondence between CFTs and TOs, so-called topological holography, potentially applicable to systems in general space-time dimensions. In Appendix C, we apply the method in the main text to the $SU(3)_{3}$ Wess-Zumino-Witten model, which is less familiar, and demonstrate a new type of fusion rules.
@@ II Bulk semionization
In this section, we present the main results of this work, the $Z_{N}$ symmetric bulk semionization of anomaly-free $Z_{N}$ two-dimensional conformal field theories (summarized as the construction of a SymTFT $\mathbf{S}^{(l)}$ from an SFC $\mathbf{F}^{(l)}$). This provides the algebraic way to relate topological symmetry or SFC and categorical symmetry or SymTFT, enabling one to calculate the fusion coefficients in general. More specifically, the derivation of $\mathbf{S}^{(l)}$ from $\mathbf{F}^{(l)}$ is new. Related approaches can be seen in [44, 47, 73, 74], but the fusion coefficients of the $Z_{N}$ graded SymTFTs are impossible (or very difficult at least) to obtain straightforwardly from their methods. We also note that the corresponding extended vertex operator algebra has been studied in the studies of monstrous moonshine[75], but their method is limited to some specific central charges, such as $c=16$ or $c=24$ (see an earlier work [76] and recent works [29, 30], for example). In the following discussion, we use the same notations as in the author’s previous works[77, 27] and note the reviews [78, 79] for earlier literature on the simple current extension.
@@
where $\times$" represents the fusion of objects, $h$ is the conformal dimension, and we do not distinguish the label and the operator itself to simplify the notations. For example, in a $Z_{2}$ model, the charge zero sector corresponds to the Neveu-Schwartz sector and the charge $1/2$ sector corresponds to the Ramond sector. Hence, the quantity $Q_{J}(\alpha)$ naturally define $Z_{N}$ analog of supercharge, fractional supercharge, or $Z_{N}$ anomaly.
The anomaly-freeness of this $Z_{N}$ theory can be defined as the condition $Q_{J}(J^{n})=0$ for all n=0,1,,N1n=0,1,...,N-1 where $J^{n}$ is the $n$ times multiplication of $J$. In other words, the $Z_{N}$ symmetry itself is not charged (or not mutually twisted) in the anomaly-free theory, and one can identify the anomaly-free theory as a fractional supersymmetric model by generalizing the arguments on the $Z_{2}$ cases. This condition can be satisfied under the following conditions,
@@
The models satisfying these conditions have been known as discrete torsion or integer simple current extension of CFTs [31, 32, 33, 34, 35, 36, 37]. For example, this condition can be realized in the $SU(N)_{N}$ Wess-Zumino-Witten (WZW) model[80, 81]. More recently, its relation to $SU(N)$ Haldane conjecture[82, 83, 84] or Lieb-Schultz-Mattis theorem has been revisited in [85, 77]. For the readers interested in the anomaly analysis of the $Z_{N}$ symmetric models and their renormalization group (RG) flows, we note several recent works [86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96] and earlier related works[97, 98, 99, 100, 101] and note recent works on noninvertible symmetry [102, 103, 104, 105, 106, 107, 108, 109, 57]. The anomaly-free $Z_{N}$ models play a central role in constructing $Z_{N}$ TOs or FQH states [110, 40, 24, 111, 112, 77, 113, 114] with close analogy and connection to quark-hadron models or fractional supersymmetric models  [^1] Following the terminologies to call fermionic (or $Z_{2}$ extended) fusion category as superfusion category, we call $Z_{N}$ extended model as fractional supersymmetric model. We also note that the $Z_{N}$ anomaly-freeness in this manuscript can be regarded as a natural extension of the notion of the anomaly-freeness in fermionic (or supersymmetric) models.. In the contexts of the FQH states and TOs, the integer simple currents define the $Z_{N}$ analog of electron operators defining the wavefunctions, and this aspect is the most fundamental in constructing the wavefunctions[38].
@@
In other words, the algebraic data of the SFC are constrained by the algebraic data of the MTC, and one can consider that a class of SFCs is determined by the combination of the Verlinde formula [5, 6, 7] and Moore-Seiberg data [68, 69]. This can be useful for the readers to understand the SFC from the knowledge of MTC or CCFT. We also note that this ring isomorphism is a fundamental structure introducing the Deligne product as a reduction from tensor products of chiral and antichiral MTCs, $\mathbf{M}\otimes\overline{\mathbf{M}}$, to the SFC $\mathbf{B}=\mathbf{M}\boxtimes\overline{\mathbf{M}}$ where $\mathbf{M}$ represents a MTC. We primarily discuss the tensor product $\otimes$, which realizes the extensions and their algebraic structure, and the fusion product $\times$ in subsequent discussions.
The simple current extended bulk CFT $\mathbf{F}$ [31, 32, 33, 34, 35, 36, 37] can be obtained by introducing the parity shift operation [129] (or recursive multiplications of $J$ and $\overline{J}$) to the bosonic theory $\mathbf{B}$. Alternatively, the simple current extension of CCFT can be realized by decomposing the $Z_{N}$ invariant sector as $\theta_{a,0}=\sum_{p=0}^{N-1}\phi_{a,p}/\sqrt{N}$ where the extended object $\phi_{a,p}$ is the $Z_{N}$ generalization of chiral semion[23, 130, 77, 27]. Because of the introduction of this extended object, the extended CCFT does not correspond to an MTC. Under the operator-state correspondence[27], the bulk operators can be specified as,
@@
where $Q$ is the $Z_{N}$ charge of the sectors, or fractional supercharge of the model. By restricting our attention to the uncharged sector with $Q=0$, the partition function completely reproduces the existing ones in the literature[31, 32, 33, 34, 35, 36, 37]. In other words, we introduced the sectors $Q\neq 0$ which can be regarded as a generalized version of the Ramond sector, and expressed the corresponding bulk theory as $\mathbf{F}$. We also note that the anomaly-free condition and the resultant partition function have been rarely discussed in the literature on category theories.
@@
As one can see from the above expressions, the summation has been taken corresponding to the addition of the anomaly-free particle $J\overline{J}^{N-1}$, and this can be interpreted as a natural extension of the boson condensation. Whereas the usual anyon condensation results in ring isomorphism or ring homomorphism, the above condensation results in a subalgebra of $\mathbf{F}$. Because of the anomaly-free condition, the total $Z_{N}$ parity and charge are well-defined after the bulk semionization. Hence, one can relate the bosonic or $Z_{N}$ orbifolded modular invariants to the Lagrangian subalgebra of $\mathbf{S}$ [145, 26, 77] based on them by taking the parity or charge 0 sector. The constraint from the charge and parity can be regarded as a natural candidate of minimal constraints with a connection to the modular invariants, and this is analogous to the Lagrangian algebra satisfying the maximality condition. More phenomenologically, the condensation should stabilize the system, and these stabilized edge modes are known as protected edge modes. The correspondence between protected edge modes and a modular invariant has been studied in earlier literature [71, 72], and their relation to condensation has been clarified in [145]. One can also see the analogy between this reduction for the modular invariant from the total partition function $\sum_{Q}Z_{Q}$ and dual models in studies on strong interactions[146, 26, 77]. The modular invariant sector, such as $Z_{0}$, corresponds to the low momentum modes in the dual model, and, by excluding (or gapping out) modular noninvariant parts corresponding to high momentum modes in the dual models, one will obtain a well-organized part of the theory corresponding to the Lagrangian subalgebra. To emphasize the role of subalgebra (or gapping operation) as a variant of anyon condensation, we use Lagrangian subalgebra instead of “Lagrangian algebra" in the existing literature.
@@ II.1 Lagrangian (sub)algebra
One can straightforwardly check their consistencies by applying the fusion rule of $\mathbf{F}^{(0)}$ and changing the basis, because $\mathbf{S}^{(0)}$ is a subalgebra of $\mathbf{F}^{(0)}$. This simplicity and rigorousness are in sharp contrast with other existing frameworks. For example, one can check the consistency of the new identity operator $\mathbf{I}$ as follows,
@@ III Topological holography and fractional supersymmetry
where we have used the same notations in Sec.II. This is the algebraic representation of topological holography[42, 43, 44, 45, 46, 47] relating a $2+1$ dimensional TO denoted by $\{\Psi\}$ and a CCFT denoted by $\{\phi\}$. As an earlier literature in mathematics, we note [49] with detailed interpretations of the Moore-Seiberg data[68, 69]. We also note that the $Z_{N}$ symmetric Cardy states in [163, 164, 165, 166, 77, 44] which has a close connection to the bulk topological degeneracies can be labeled by the chiral semions because of the close connection to CCFTs and boundary CFTs (BCFTs)[63, 64, 167, 168, 169, 170] or the Li-Haldane conjecture in condensed matter[65, 66, 67]. In existing literature, the extended algebraic structure $\mathbf{S}$ has been obtained heuristically from the algebraic data of $\mathbf{M}$ in some particular cases. Our arguments systematically provide $\mathbf{S}$ as a subalgebra of $\mathbf{F}$, and this is in sharp contrast to the other existing arguments. We stress that even the categorical studies on the extended theory $\mathbf{F}$ are still under development.